OpenStudy (anonymous):

describe the vertical asymptotes and holes for the graph of y=(x-4)/(x^2+3x+2)

OpenStudy (anonymous):

@Luigi0210

OpenStudy (luigi0210):

Do you know how to find vertical asymptotes and holes?

OpenStudy (anonymous):

Not realy

OpenStudy (anonymous):

@Luigi0210

OpenStudy (luigi0210):

Well okay, can you factor the denominator?

OpenStudy (anonymous):

(x-4)/(x+2)(x+1)

OpenStudy (luigi0210):

And just by that, we see it has no holes. So now for asymptotes, set each factor in the denominator equal to 0 and solve

OpenStudy (anonymous):

-4/2=-2

OpenStudy (luigi0210):

That's one, now find the other

OpenStudy (anonymous):

But what do I plug in because I already pluged 0 in

OpenStudy (luigi0210):

No, I mean solve them like this: \[\LARGE x+1=0\] and \[\LARGE x+2=0\]

OpenStudy (anonymous):

confused...

OpenStudy (luigi0210):

When looking for asymptotes, you have to set it equal to zero and solve: \[\LARGE 0=x^2+3x+2\]

OpenStudy (anonymous):

So -1 @Luigi0210

OpenStudy (luigi0210):

Yea, and -2

OpenStudy (anonymous):

Thanks

OpenStudy (luigi0210):

Your graph will look something like this, if you're graphing it.